# Multimodal 3D Reasoning Segmentation with Complex Scenes

### v3

@@ 5.1 Experimental Settings
We conduct experiments on one NVIDIA V100 GPU (32 GB memory) and train the network for 100 epochs with a batch size of 1, with a total training time of around 1 day. We use the Adam [23] optimizer with the initial learning rate of $1\times 10^{-4}$. We adopt LLaMA-7B [38] as our multimodal LLM backbone and LoRA [15] to perform efficient fine-tuning. All experiments are conducted on the proposed ReasonSeg3D dataset.
Table 2: Benchmarking on the ReasonSeg3D validation set with evaluation metrics cIoU and gIoU. Best in bold, second underlined.
         Method          Venue          cIoU          gIoU
         3D Segmentation Methods
         OpenScene [32]          CVPR 23          7.69          9.52
         PLA [9]          CVPR 23          10.76          10.27
         RegionPLC [48]          CVPR 24          10.82          11.06
         3D-STMN [43]          AAAI 24          19.37          20.70
         RefMask3D [12]          MM 24          20.53          22.81
         MDIN [42]          MM 24          23.04          23.49
         3D General Scene Understanding Methods
         3D-LLM [14]          NeurIPS 23          21.36          22.49
         3D-VisTA [58]          ICCV 23          23.95          24.77
         3D Reasoning Segmentation Methods
         Reason3D [16]          3DV 25          25.12          25.90
         MORE3D (Ours)          -          30.19          32.01
Figure 5: Segmentation visualization over the ReasonSeg3D validation set. Each case presents a user input question, the corresponding input point cloud, the ground-truth segmentation, and the prediction by the proposed MORE3D. Best viewed in color and zoom-in. Green indicates chairs, and pink indicates tables. Best viewed in color.
5.2 Benchmarking with Existing Methods
We benchmark MORE3D with state-of-the-art methods, including 3D segmentation [32, 9, 48, 43, 12, 42], general scene understanding [14, 58], and reasoning segmentation [16] approaches, on the ReasonSeg3D validation set. All compared methods are trained and evaluated on the same ReasonSeg3D to ensure a fair comparison. As Table 2 shows, our method achieves superior reasoning segmentation performance across all evaluation metrics. Specifically, OpenScene [32], PLA [9], and RegionPLC [48] are designed to comprehend input words or phrases based on CLIP-based language models. They lack the ability to reason user intention from input questions and to correctly generate answers targeting multiple corresponding objects. Consequently, the predicted 3D segmentation masks by these methods cannot be aligned correctly with ground-truth 3D segmentation masks, resulting in low cIoU and gIoU. Moreover, 3D-STMN [43] and RefMask3D [12] are tailored for single-object segmentation and therefore cannot handle multiple objects, while MDIN [42] exhibits weaker reasoning ability for interpreting implicit human intention. Additionally, 3D-LLM [14] and 3D-VisTA [58], designed for general 3D tasks, are re-implemented on our proposed 3D reasoning segmentation task for comparison. Furthermore, Reason3D [16], as a 3D reasoning segmentation approach, is limited to single-object settings and struggles with multi-object reasoning segmentation. In contrast, MORE3D can accurately reason the implicit intention of user questions, predicting 3D segmentation masks that are more consistent with the ground truth.
Qualitative Results.
5.3 Ablation Study
We conduct ablation studies on the ReasonSeg3D validation set to evaluate our designs. Specifically, we examine MORE3D from the aspect of the approach for object-specific embedding extraction, the way of prediction, the point cloud decoding approach, as well as the loss design.
Embedding Extraction.
We examine the effectiveness of the designed multi-seg index list $I_{seg}$ for the object-specific embedding extraction in MORE3D. As Table 6 shows, using the multi-seg index list $I_{seg}$ can achieve better performance than adopting the random selection approach. This is largely because the LLM’s embeddings corresponding to multiple objects are extracted precisely with the help of $I_{seg}$. As a comparison, the random selection could extract embeddings unrelated to target objects like “table" or “chairs", and it could also extract embeddings belonging to irrelevant non-object words, such as “around" and “suitable". This explains why using the multi-seg index list $I_{seg}$ is more effective for extracting object-specific point cloud embeddings corresponding to multiple target objects.
Table 3: Ablation study on object-specific embedding extraction approach.
Index Approach cIoU gIoU
1 Random 20.30 21.23
2 with $I_{seg}$ 30.19 32.01
@@ 5.3 Ablation Study
Table 5: Ablation study on the approach of point cloud decoding operation in the 3D Decoder.
Index Operation cIoU gIoU
1 Addition 28.26 31.01
2 Concatenation 29.42 31.17
3 Dot Product 30.19 32.01
@@
We examine the effectiveness of different prediction approaches including unified and separated approaches for the prediction of 3D segmentation masks and classification. Table 6 shows experimental results. For the unified approach, the 3D segmentation masks and classification results are predicted by a single network instead of two separate network branches. This results in lower performance due to the increased learning burden when a single network branch must learn to perform segmentation and classification simultaneously. In contrast, the separate approach as illustrated in the bottom right of Figure 3 assigns each task to an independent network branch to learn. This allows each branch to focus on its respective task which lowers the learning burden and improves performance clearly.
Decoding Operation.
We examine how different point cloud decoding operations affect 3D reasoning segmentation. Besides the dot product for refined queries and features, we explore addition and concatenation to investigate whether richer point cloud information improves training. As Table 6 shows, the Addition and the Concatenation operations achieve lower performance compared with the Dot Product operation since they either add or concatenate the embeddings of each object with the per-point features of the global scene, introducing abundant features of irrelevant objects and leading to inferior performance. Differently, the Dot Product operation selectively extracts only the relevant features from the per-point features of the global scene by identifying those highly correlated with the object embeddings. This leads to better performance thanks to its extracted object-specific features.

### v4

@@ Introduction
We conduct experiments on one NVIDIA V100 GPU (32 GB memory) and train the network for 100 epochs with a batch size of 1, with a total training time of around 1 day. We use the Adam (Kingma 2015) optimizer with the initial learning rate of $1\times 10^{-4}$. We adopt LLaMA-7B (Touvron et al. 2023) as our multimodal LLM backbone and LoRA (Hu et al. 2021) to perform efficient fine-tuning. All experiments are conducted on the proposed ReasonSeg3D dataset.
Figure 4: Segmentation visualization over the ReasonSeg3D validation set. Each case presents a user input question, the corresponding input point cloud, the ground-truth segmentation, and the prediction by the proposed MORE3D. Green indicates chairs, and pink indicates tables. Underlined text denotes 3D spatial relations. Best viewed in color and zoom-in.
@@
Benchmarking with Existing Methods
We benchmark MORE3D with state-of-the-art methods, including 3D segmentation (Peng et al. 2023; Ding et al. 2023; Yang et al. 2024b; Wu et al. 2024c; He and Ding 2024; Wu et al. 2024b), general scene understanding (Hong et al. 2023; Zhu et al. 2023), and reasoning segmentation (Huang et al. 2025) approaches, on the ReasonSeg3D validation set. All compared methods are trained and evaluated on ReasonSeg3D to ensure a fair comparison. As Table 2 shows, our method achieves superior performance across all evaluation metrics. Specifically, OpenScene (Peng et al. 2023), PLA (Ding et al. 2023), and RegionPLC (Yang et al. 2024b) are designed to comprehend input words or phrases based on CLIP-based language models. They lack the ability to reason user intention from input questions and to correctly generate answers targeting multiple corresponding objects. Consequently, the predicted 3D segmentation masks by these methods cannot be aligned correctly with ground-truth 3D segmentation masks, resulting in low cIoU and gIoU. Moreover, 3D-STMN (Wu et al. 2024c) and RefMask3D (He and Ding 2024) are tailored for single-object segmentation and therefore cannot handle multiple objects, while MDIN (Wu et al. 2024b) exhibits weaker reasoning ability for interpreting implicit human intention. Additionally, 3D-LLM (Hong et al. 2023) and 3D-VisTA (Zhu et al. 2023), designed for general 3D tasks, are re-implemented on our proposed 3D reasoning segmentation task for comparison. Furthermore, Reason3D (Huang et al. 2025), as a 3D reasoning segmentation approach, is limited to single-object settings and struggles with multi-object reasoning segmentation. In contrast, MORE3D can accurately reason the implicit intention of user questions, predicting 3D segmentation masks that are more consistent with the ground truth.
Qualitative Results.
Ablation Study
We conduct ablation studies on the ReasonSeg3D validation set to evaluate our designs. Specifically, we examine MORE3D with the approach for the way of prediction, and the loss design.
@@
We examine the effectiveness of different prediction approaches including unified and separated approaches for the prediction of 3D segmentation masks and classification. Table 3 shows experimental results. For the unified approach, the 3D segmentation masks and classification results are predicted by a single network instead of two separate network branches. This results in lower performance due to the increased learning burden when a single network must learn to perform segmentation and classification simultaneously. In contrast, the separate approach as illustrated in Figure 2 assigns each task to an independent network branch to learn. This allows each branch to focus on its respective task, which lowers the learning burden and improves performance clearly.